OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 2x+3y=5 and contains the point (-2,1).

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

the slope of 2x +3y = 5 is -2/3 so y - y1 = -2/3(x - x1) where y1 = 1 and x1 = -2

OpenStudy (anonymous):

so y -1 = -2/3x + 4/3 times everything by 3 3y - 3 = -2x +4 3y +2x = 7

OpenStudy (anonymous):

no they look the same but are different

OpenStudy (anonymous):

yeah i noticed that, sorry

OpenStudy (anonymous):

oh its okay lol. but your answer doesnt match any of my answers? :(

OpenStudy (anonymous):

yeah i messed up on my addition

OpenStudy (anonymous):

it should be 3y - 3 = -2x - 4 so the answer is 3y +2x = -1 answer D

OpenStudy (anonymous):

thank you very much! :)

OpenStudy (anonymous):

you're welcome